jadeleavitt:

Solve the system y = -x + 7 and y = -0.5(x - 3)^2 + 8

Imagine:

https://questioncove.com/study#/updates/5fc295d057639d2d6bb07c00 (FINDING THE X TUTORIAL)

Timmyspu:

Ok so first off do you have a algebraic calculator

jadeleavitt:

yes

jhonyy9:

do you need help ?

jadeleavitt:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 do you need help ? \(\color{#0cbb34}{\text{End of Quote}}\) yes please

jhonyy9:

do you know the substitution method ?

jhonyy9:

or elimination method ?

jhonyy9:

what you wan or need use here ?

Mercury:

old question, but will respond so this can be closed y = -x + 7 and y = -0.5(x - 3)^2 + 8 you can set them equal to each other to get -x + 7 = -0.5(x-3)^2 + 8 you can expand the right side using FOIL, then combine the constant terms -x + 7 = -0.5x^2 + 3x + 3.5 from there, you can add x - 7 to both sides, to get 0 on one side 0.5x^2 - 4x + 3.5 = 0 and use the quadratic formula to get your possible x-values. from there, plug x into the equation to find the corresponding y-values.